Given that p+q+r=0. two of them are equal in magnitude. the magnitude of third vector is √2  times that of the other two. which are the angles between these two vectors?
90, 135, 135
45, 45, 135
30, 60, 90
45, 90, 135
plz answer fast..

Grade:11

1 Answers

Mohsin syed
5 years ago
Given:P+Q+R=0
SO,
P+Q=-R
Given P=Q
P²+p²+2p²Cos∆=-q²
2p²+2p²cos∆=(p√2)²
2p²Cos∆=2p²-2p²
Cos∆=0
∆= 90 degree= angle btw p & q
 
Now btw p and R
P² +(p √2)²+2(p)(p√2)cos ∆= -q²
P²+ 2p² +2p²√2cos∆=p²
2p²√2cos∆=-2p²
Cos∆= -2p²/2p²√2
Cos∆ =-1/√2
∆= 135 degree 
Hence angles are 90,135 q35